OpenStudy (anonymous):

Find the horizontal asympote, if any, of the rational function. f(x)=5/x^2+1

myininaya (myininaya):

y=0 is a horizontal asy

OpenStudy (anonymous):

y=0 is a vertical asymptote. There's a horizontal one at y=1

myininaya (myininaya):

just divide x^2 on both top and bottom 5/x^2 goes to 0 as x goes to infinity

myininaya (myininaya):

\[y=\frac{5}{x^2+1}?\]

OpenStudy (anonymous):

\[5/x ^{2}+1\] is what I read, which explains the discrepancy

myininaya (myininaya):

did you mean y=5/(x^2+1) or y=5/x^2+1 i just assumed he doesnt write parenthesis like the last problem

myininaya (myininaya):

you could be right

OpenStudy (anonymous):

If it reads as myin read it, there are no asymptotes. If it reads as I read it, there's a vert at x=0 and a horizont at y=1

OpenStudy (anonymous):

the original problem doesn't have parenthesis its written f(x)=5/x^2+1

OpenStudy (anonymous):

myininaya wrote it correct

OpenStudy (anonymous):

so the answer is "none"

myininaya (myininaya):

\[y=\frac{5}{x^2+1}\] you can write that as 5/(x^2+1) but if you write it as y=5/x^2+1, you mean\[y=\frac{5}{x^2}+1\] so we have \[y=\frac{5}{x^2+1}\] we have y=0 is a horizontal asymptote

OpenStudy (anonymous):

yes, the way that you wrote it the first time is the correct way to write the problem.

myininaya (myininaya):

lol